Order: ​DTIC-Dome (dacarbazine) 100​ mg/m2 IV daily for 5 days. The label on the vial states 100 mg to reconstitute with 9.9 mL of sterile water for​ injection, yielding a strength of 100​ mg/10 mL, further dilute in 100 mL of​ D5W, and infuse over 60 minutes. Calculate the rate of flow in milliliters per hour for a patient who weighs 70 kg and is 180 cm tall.

A.

250​ mL/hr

B.

304​ mL/hr

C.

299​ mL/hr

D.

206​ mL/hr
Answer #1

ORDER = DTIC- DOME

100 mg/ m2 iv For 5 days

100 MG DILUTED IN 10 ML = 10 ML

FURTHER DILUTION IN 100 ML D5W= 100 ML

TOTAL= 100+ 10 = 110 ml

patients weight = 70 kg

patients height = 180 cm tall

calculating body surface area=   (Mosteller)5 =height(cm) x weight(kg) 3600

√3.5 = 1.87

1.87 IS THE BSA

total amount the patient will receive is 100 X 1.87= 187 MG

UNKNOWN AMOUNT= X

DESIRED AMOUNT = D

ON HAND AMOUNT = H

VOLUME OF DILUTION = Q

X= D/H*Q

X= 187MG / 100 MG *11O ML

X= 205.7ML (VOLUME NEEDED)

RATE OF FLOW= VOLUME IN (ML) / TIME IN (HOUR)

= 205.7 ML / 1 HR (60 MIN GIVEN)

= 205.7 ML/HR

APPROX 206 ML/ HR

OPTION (D) IS THE CORRECT ANSWER THAT IS( 206 ML/ HR)